OpenStudy (anonymous):

I'm so confused on this problem can someone solve it and show steps I need to see where I'm going wrong. 5,000=P(1+ 00.45/12) ^12*15

OpenStudy (anonymous):

5,000=p(1+0.045/12)^12*15

OpenStudy (neer2890):

\[5000=p(1.45)^{\frac{ 4 }{ 5 }}=p*1.35\] \[p=\frac{ 5000 }{ 1.35 }\]

OpenStudy (anonymous):

@neer2890 Not sure I understand why you combined 12*15 to raise the group to a power of 4/5. Wouldn't the group be raised to a power of 12 first then the (*15) be applied? Here is my interpretation of this equation - \[5000=p(1+(0.045/12))^{12}*15\] Because the problem was written 2 different ways, (0.045 and 0.45) the actual math may be different, but, unless I missed something liked implied grouping of exponents, I simply don't see this equation as you do. If I am incorrect, please let me know, as I don't want to give out bad info. The way I understand orders of operations, I would not raise to a power of 12/15 unless the original equation was written -\[5000=p(1+0.045/12)^{12*15}\]. Also, what happened to dividing the 0.45 (or 0.045) by 12? I am not trying to be critical, just want to understand how you got to your conclusion.

OpenStudy (amistre64):

this looks like a compound interest formula\[A=P(1+r/n)^{nt}\] \[\frac{A}{(1+r/n)^{nt}}=P\]

OpenStudy (neer2890):

i think i misunderstood the question. @Felix820 \[50000=p(1+\frac{ 0.045 }{ 12 })^{12*15}\] \[5000=p(1.0037)^{180}=p*1.96\] \[p=\frac{ 5000 }{ 1.96 }\]

OpenStudy (anonymous):

@neer2890 Didn't mean anything by it. I'm brand new here, been outa school for way too long to reveal, and just wanted to make sure I wasn't brain dead or worse! LOL! Math WAS the one subject I did well in oh so many years ago. Stumbled onto this site quite by accident. I'm having fun testing my skills and recall! Thanks for the msg.

OpenStudy (neer2890):

nvm buddy... and thanks for correcting me...!

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

i was 20 years out before i went back to college. just wrapped up a BA in mathematics this May.

OpenStudy (anonymous):

@amistre64 WTG!

OpenStudy (anonymous):

@Felix820 still there?